OpenStudy (anonymous):

Ca somone show me how to solve Sin-1 (- square root 3/2)

OpenStudy (anonymous):

someone*

OpenStudy (anonymous):

can't really show how to solve but you need to know your unit circle the answer is 7pi/3 and 11pi/3

OpenStudy (psymon):

sin^-1 and sin are inverse. They basically flip-flop values around. Normally we have: sinx = some number. But what sin^-1 does is say sin^-1(some number) = x You can always move these around in this way. So when your problem says sin^-1 (-sqrt(3)/2) = x, I can rewrite that as sinx = -sqrt(3)/2 So if you dont know the answer, you can always look on the unit circle and check what angle of sin has a value of -sqrt(3)/2 That kind of make sense?

OpenStudy (anonymous):

I have the unit circle and the book says its - pi over 3 , but how?

OpenStudy (psymon):

Well, I can explain that, I just want to make sure you get what I was trying to say above.

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

sqrt3/2 for sin is pi/3 but because it was negative the answer is in the 3rd and 4th quadrants

OpenStudy (psymon):

Kinda make sense? If so, Ill explain the answer.

OpenStudy (anonymous):

right costanza, its only in the 3rd and 4th quarters, but i guess what im trying to find out if its in the 3rd and 4th q, how or where does it = to pi/3

OpenStudy (psymon):

I can explain that, but only if you let me. Otherwise ill let costanza do it. Im guessing what I was trying to say didnt make sense then?

OpenStudy (anonymous):

no it did not make since sorry.. (lost) , thats fine please explain psymon..

OpenStudy (psymon):

Basically, if you ever have sin^-1(y) = x, you can rewrite it as sinx = y. So just examples. sin(pi/2) = 1 becomes sin^-1(1) = pi/2 cos(pi) = -1 becomes cos^-1(-1) = pi As for yours sin^-1(-sqrt(3)/2) = x sinx = -sqrt(3)/2

OpenStudy (psymon):

Not sure if you catch the pattern or not, though.

OpenStudy (anonymous):

o i c, got it! thanks for explaining!! =)

OpenStudy (psymon):

Yep. Now for the answer! inverse sin MUST MUST MUST be in between -pi/2 and pi/2. So even though sin^-1(-sqrt(3)/2) can be 4pi/3, that is in the 3rd quadrant, which is NOT in between -pi/2 and pi/2. So basically, sin^-1 only has answers in quadrant 1 and 4, nowhere else. Now the two values for which sin is normally -sqrt(3)/2 are 4pi/3 and 5pi/3. Well, 5pi/3 is in the correct quadrant, but it is not in between -pi/2 and pi/2. So instead of calling it 5pi/3, we call it by what the same angleof it is negative, which is -pi/3